Calculate the amount of concentrated 11.6 M HCl that is necessary to prepare 500.00 mL of a 1.00 M HCl solution? Report your answer to the correct number of significant figures.
Answer #1

use dilution formula
M1*V1 = M2*V2
1---> is for stock solution
2---> is for diluted solution

Given:
M1 = 11.6 M
M2 = 1.00 M
V2 = 500.00 mL

use:
M1*V1 = M2*V2
V1 = (M2 * V2) / M1
V1 = (1*500)/11.6
V1 = 43.1 mL
Answer: 43.1 mL